OpenStudy (anonymous):

Which list of ordered pairs represents solutions to x + y = 2 ? (-4, 6), (0, 2), (4, 2) (-4, -6), (0, 2), (4, 2) (-4, 6), (0, 2), (4, -2)

OpenStudy (anonymous):

try plugging in those choices.... which one woks for all the given points?

OpenStudy (anonymous):

if i could do math i would.

OpenStudy (anonymous):

C

OpenStudy (anonymous):

lol... you can at least add, right?

OpenStudy (anonymous):

dont be dumb of course i can

OpenStudy (anonymous):

my bad.... sorry....

OpenStudy (anonymous):

so is it C?

OpenStudy (anonymous):

(-4, 6), (0, 2), (4, -2) let me be dumb again for a minute.... -4 + 6 = 0 + 2 = 4 + -2 = now what do you think?

OpenStudy (anonymous):

i think i was correct